OpenStudy (anonymous):

log(base4)(x+3)+log(base4)(2-x)=1

OpenStudy (debbieg):

Solve for x? Use rules of logs to combine on left side. \[\log_{4}(x+3) +\log_{4}(2-x)=\log_{4}[(x+3) (2-x)]=\log_{4}[(x^2-x+6)]\] So now you have \[log_{4}[(x^2-x+6)]=1\] Convert to exponential form:\[4^{1}=x^2-x+6\]\[x^2-x+6=4\]\[x^2-x+2=0\] Now solve this quadratic equation. Beware of extraneous solutions.

OpenStudy (anonymous):

\[\log_{4} \left( x+3 \right)+\log_{4} \left( 2-x \right)=1\] \[\log_{4} \left( x+3 \right)\left( 2-x \right)=1\] \[4^{1}=\left( x+3 \right)\left( 2-x \right)\] now you can solve as above

OpenStudy (anonymous):

thank you both! :)

OpenStudy (anonymous):

\[4=2x-x ^{2}+6-3x orx ^{2}+x-2=0\] \[x ^{2}+2x-x-2=0,\] x(x+2)-1(x+2)=0 (x+2)(x-1)=0 either x=-2 or x=1

OpenStudy (anonymous):

yw